# Quantum symmetries of noncommutative tori

### v1

@@ 3.2. Actions of Haagerup-Izumi on noncommutative 2-tori
    $\mathcal{H}^{*}$ contains a copy of $\text{Hib}(\widehat{G})$, and each $\gamma\in\widehat{G}$ acts trivially as a functor on $\mathcal{M}$
@@ 3.3. Action of adjoint subcategory of $E_{8}$ quantum subgroup on a noncommutative 3-torus
With the natural ordering $m_{1}=\mathbbm{1},m_{2}=A,m_{3}=B$ and $m_{4}=\tau$, the fusion matrix for $X$ (and hence the connecting diagram for $\iota_{n}$) is
@@
$\Theta$ is non-degenerate, hence $A_{\Theta}$ is a simple AT algebra with unique trace.
@@ 3.4. Action of $\text{PSU}(2)_{15}$ on a noncommutative 4-torus
Let $\mathcal{F}:=PSU(2)_{15}$, where we use notations and conventions as described in Section 3.1. This fusion category is rank $8$, and we consider $\mathcal{M}:=\mathcal{F}$ as a right $\mathcal{F}$-module category, and we set $m=\sum_{a}Y_{a}$. Note $\mathcal{F}^{*}_{\mathcal{M}}\cong\mathcal{F}$. Then set $X=Y^{\otimes 4}_{1}$. The triple $(\mathcal{M},m,X)$ satisfies the hypothesis of Theorem 2.13, and thus we obtain an action of $\mathcal{F}$ on the simple AT-algebra $B$ with unique trace and $K_{1}(B)\cong\mathbbm{Z}^{8}$.
@@
Acknowledgements. This work was initiated during the authors visits to the Hausdorff Research Institute in Bonn in August 2022 for the follow up workshop to the earlier programme on von Neumann algebras. DEE is supported by a Leverhulme Emeritus Fellowship. Part of the work was conducted when he visited Dublin Institute for Advanced Studies, the University of New South Wales and Harvard University. He would like to thank Denjoe O’Connor, Pinhas Grossman (and the Australian Research Council Discovery Project DP170103265) and Arthur Jaffe for their generous hospitality and stimulating environments. CJ is supported by NSF Grant DMS-2247202. We are grateful to George Elliott, Sergio Girón Pacheco, and Andrew Schopieray for valuable comments.

### v3

@@ 3.2. Actions of Haagerup-Izumi on noncommutative 2-tori
    $\mathcal{H}^{*}$ contains a copy of $\text{Hilb}(\widehat{G})$, and each $\gamma\in\widehat{G}$ acts trivially as a functor on $\mathcal{M}$
@@ 3.3. Action of adjoint subcategory of $E_{8}$ quantum subgroup on a noncommutative 3-torus
With the natural ordering $m_{1}=\mathbbm{1},m_{2}=A,m_{3}=B$ and $m_{4}=\tau$, the fusion matrix for $X\otimes{\cdot}$ (and hence the connecting diagram for $\iota_{n}$) is
@@
[…]
@@ 3.4. Action of $\text{PSU}(2)_{15}$ on a noncommutative 4-torus
[…]
@@
Acknowledgements. This work was initiated during the authors visits to the Hausdorff Research Institute in Bonn in August 2022 for the follow up workshop to the earlier programme on von Neumann algebras. DEE is supported by an Emeritus Fellowship from the Leverhulme Trust. Part of the work was conducted when he visited Dublin Institute for Advanced Studies, the University of New South Wales and Harvard University. He would like to thank Denjoe O’Connor, Pinhas Grossman (and the Australian Research Council Discovery Project DP170103265) and Arthur Jaffe for their generous hospitality and stimulating environments. CJ is supported by NSF Grant DMS-2247202. We are grateful to Alexander Betz, George Elliott, Sergio Girón Pacheco, and Andrew Schopieray for valuable comments. We also thank Kan Kitamura for pointing out an error in a previous version of the manuscript.